OpenStudy (cookie_2046):

Simplify -4(x-6)+3(a-7) A. -4x+3a-45 B. -4x+3a+45 C. -4x+3a-3 D.-4x+3a+3 Idk this one

OpenStudy (cookie_2046):

Gonna guess.. I think maybe b? Not so sure..

OpenStudy (anonymous):

what is the question

OpenStudy (cookie_2046):

Simplify -4(x-6)+3(a-7)

OpenStudy (anonymous):

@KaylaRdz0405 she will be more than glade to help u she is smart and she knows this one

OpenStudy (cookie_2046):

Alright (:

OpenStudy (anonymous):

D.

OpenStudy (cookie_2046):

thank you!

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

you got it right?

OpenStudy (cookie_2046):

idk yet i need to submit my quiz and i still have 1 more question to complete

OpenStudy (anonymous):

k hold on

OpenStudy (michele_laino):

If I apply the distributive property of multiplication over addition, I can write this: \[\Large \begin{gathered} - 4\left( {x - 6} \right) + 3\left( {a - 7} \right) = \hfill \\ \hfill \\ = - 4x + 24 + 3a - 21 = ...? \hfill \\ \end{gathered} \]

OpenStudy (cookie_2046):

-4(x-2)-5(y+2)+(-3)(8-6b) A. -4x-5y+18b-6 B.-4x-5y+18b+6 C.-4x-5y-18b-22 D. -4x-5y+18b-26 Do u know this last one?

OpenStudy (cookie_2046):

Simplify -4(x-2)-5(y+2)+(-3)(8-6b)

OpenStudy (michele_laino):

again, if we apply the distributive property of multiplication over addition, we can write this: \[\Large \begin{gathered} - 4\left( {x - 2} \right) - 5\left( {y + 2} \right) + \left( { - 3} \right)\left( {8 - 6b} \right) = \hfill \\ \hfill \\ = - 4x + 8 - 5y - 10 - 24 + 18b = ...? \hfill \\ \end{gathered} \]

OpenStudy (cookie_2046):

D is answer correct?

OpenStudy (anonymous):

:D

OpenStudy (michele_laino):

yes! correct!

OpenStudy (anonymous):

:)